# Efficient Bayesian Updates for Deep Learning via Laplace Approximations

### v2

@@ 5 Use Case: Deep Active Learning
Figure 7: Accuracy improvement curves showing the difference in accuracy between the respective selection strategy and random instance selection.
@@
For this use case, we choose the CIFAR-10 dataset with a label budget of $B=200$, an acquisition size of $b=10$ instances, and the same DNN architecture as in our previous experiments (cf. Section 4). Figure 7 shows the accuracy improvement curves relative to a random instance selection. A more detailed experimental setup and learning curves reporting absolute values are available in Appendix F. In general, the resulting curves confirm our hypothesis. All query strategies using our updates outperform the respective top-$b$ selection strategies. Furthermore, combining our update with the batch selection strategy Badge also results in improved accuracy. This indicates that our idea of selecting a single instance and updating the DNN leads to a more effective selection than using the $k$-means algorithm as proposed in Badge.
6 Discussion and Conclusion
In this article, we proposed a second-order Bayesian update method for DNNs based on a last-layer LAs. It provides an efficient alternative to retraining for scenarios where data arrives sequentially and immediate updates are beneficial (e.g., AL). A extensive experimental evaluation showed that the proposed updates are both fast and effective, consistently outperforming competitors. In an AL use case, we demonstrated that the proposed update can improve sequential AL selection strategies such as Margin and BALD, as well as the popular batch selection strategy Badge.
Figure 8: SNGP with our update. See Fig. 1 for the legend.
While our proposed update has been shown to work well in the context of AL, further research is needed to explore its effectiveness in other scenarios. Specifically, when working with limited computational resources, our update can offer a efficient way to improve performance [40]. In this context, we also plan to investigate the hyperparameter $\gamma$ in greater depth to adapt it during the learning process. Additionally, while our updates are effective for vanilla DNNs combined with a last-layer LA, further validation is needed for more recent models. As an initial step, we took the 2D example from Section 1 and combined our update with SNGP [30], a model designed for accurate uncertainty modeling. Figure 8 shows that the resulting update is both effective and preserves SNGP’s uncertainty modeling capabilities.
Finally, we plan to utilize our updates to enhance or develop one-step-look-ahead selection strategies [44] in deep AL. As these selection strategies are based on decision theoretic principles, they naturally balance explorative and exploitative instance selection, a key challenge in AL [28]. However, the integration of these strategies into deep learning is challenging. Current research either relies on MC-based updates [50], which we have shown to be ineffective, or does not consider deep learning [23]. Our updates offer a new and effective way to combine such strategies with DNNs.
@@ A.1 MC-based Bayesian Updates
MC-based Bayesian neural networks (BNNs) such as deep ensembles and MC-Dropout draw samples, or hypotheses, from an (approximate) posterior distribution $q(\boldsymbol{\omega}|\mathcal{D})$. To obtain these samples, deep ensembles train multiple randomly initialized deep neural networks (DNNs), while MC-Dropout randomly sets a portion of parameters to zero for multiple inference steps. We refer to [15] for an in-depth explanation about these techniques.
@@
$q(\boldsymbol{\omega}_{m}|\mathcal{D}^{\oplus},\mathcal{D})=\text{Cat}(m|\hat{{\boldsymbol{p}}}^{\text{upd}})$ $\propto q(\boldsymbol{\omega}_{m}|\mathcal{D})\prod_{({\boldsymbol{x}},y)\in\mathcal{D}^{\oplus}}p(y|{\boldsymbol{x}},\boldsymbol{\omega}_{m})$ (13)
$=\hat{p}_{m}\prod_{({\boldsymbol{x}},y)\in\mathcal{D}^{\oplus}}\left[\mathrm{softmax}(f^{\boldsymbol{\omega}_{m}}({\boldsymbol{x}}))\right]_{y}=\hat{z}_{m},$ (14)
which is a categorical distribution with parameters $\hat{{\boldsymbol{p}}}^{\text{upd}}=(\hat{p}^{\mathrm{upd}}_{1},\ldots,\hat{p}^{\mathrm{upd}}_{M})^{\mathrm{T}}$ that we obtain after normalizing ${\boldsymbol{\hat{z}}}=(\hat{z}_{1},\ldots,\hat{z}_{M})^{\mathrm{T}}$. Intuitively, the importance of each hypothesis is determined by its likelihood of explaining the new dataset $\mathcal{D}^{\oplus}$. This approximation $q(\boldsymbol{\omega}_{m}|\mathcal{D}^{\oplus},\mathcal{D})$ of the posterior distribution allows us to make new predictions by evaluating the predictive distribution from Eq. (1) accordingly:
@@
which is a weighted average of the sampled hypotheses. Employing the update from Eq. (14) may lead to catastrophic forgetting. Thus, to control the influence of the new dataset $\mathcal{D}^{\oplus}$, we introduce the hyperparameter $\gamma$:
$q(\boldsymbol{\omega}_{m}|\mathcal{D}^{\oplus},\mathcal{D})$ $\propto\hat{p}_{m}\left(\prod_{({\boldsymbol{x}},y)\in\mathcal{D}^{\oplus}}\left[\mathrm{softmax}(f^{\boldsymbol{\omega}_{m}}({\boldsymbol{x}}))\right]_{y}\right)^{\gamma}.$ (17)
@@ A.2 First-Order Bayesian Updates
Our proposed update method from the main text approximates the new posterior $q(\boldsymbol{\omega}|\mathcal{D},\mathcal{D}^{\oplus})$ by applying an optimization step via Gauss-Newton and estimating the new covariance. Naturally, we also introduce a first-order update, employing a first-order optimization step, leading to a less complex and faster method. This also allows us to ablate the importance of the Hessian.
Assume we have a approximate posterior distribution $q(\boldsymbol{\omega}|\mathcal{D})=\mathcal{N}(\boldsymbol{\omega}|\hat{{\boldsymbol{\mu}}},\hat{{\boldsymbol{\Sigma}}})$ from an LA and observe the new dataset $\mathcal{D}^{\oplus}$. Our goal is to update the approximate posterior distribution $q(\boldsymbol{\omega}|\mathcal{D})$ by considering it as the new prior distribution:
$q(\boldsymbol{\omega}|\mathcal{D},\mathcal{D}^{\oplus})=\mathcal{N}(\boldsymbol{\omega}|\hat{{\boldsymbol{\mu}}}^{\text{upd}},\hat{{\boldsymbol{\Sigma}}}^{\text{upd}})\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}q(\boldsymbol{\omega}|\mathcal{D})\prod_{({\boldsymbol{x}},y)\in\mathcal{D}^{\oplus}}p(y|{\boldsymbol{x}},\boldsymbol{\omega}).$ (20)
@@ Appendix B Hyperparameter Ablation
For all introduced Bayesian update methods, the hyperparameter $\gamma$ controls the influence of the new dataset $\mathcal{D}^{\oplus}$ on the posterior distribution $q(\boldsymbol{\omega}|\mathcal{D})$. In Section 4, we conducted a small ablation study, similar to a hyperparameter search, to determine an appropriate value for $\gamma$. We intentionally did not search a optimal value of $\gamma$ for every dataset since extensive hyperparameter search for update methods is impractical in an online setting [8]. Using the same setup as in Section 4, we consider the CIFAR-10 and DBPedia datasets. Our baseline DNN is trained on a randomly sampled initial dataset $\mathcal{D}$ of 50 instances. We then update and retrain the DNN with varying sizes of the new dataset $|\mathcal{D}^{\oplus}|\in\{1,\dots,10\}$.
@@
Accuracy
Size of $\mathcal{D}^{\oplus}$
Refer to captionSize of $\mathcal{D}^{\oplus}$
Refer to captionSize of $\mathcal{D}^{\oplus}$
Refer to captionSize of $\mathcal{D}^{\oplus}$
@@ Appendix C Multi-class Last-Layer Update
Spiegelhalter and Lauritzen [49] presented a more efficient approximation in which the class-wise parameter vectors are arranged column-wisely as a matrix. Their joint distribution is then modeled through a matrix normal distribution:
@@
Liu et al. [29] presented an even faster approximation and showed its effectiveness in combination with an LA for supervised learning. Therefore, we employ this approximation in the multi-class setting. The idea is to determine an upper-bound covariance matrix shared by all class-wise parameter vectors in their respective multivariate normal distribution, which is then defined for class $y\in\mathcal{Y}$ as:
@@
Analog to the updates for binary classification in Eq. (9) and Eq.(10), we implement the updates of the mean parameter vector for class $y\in\mathcal{Y}$ and the covariance matrix $\hat{{\boldsymbol{\Sigma}}}\in\mathbb{R}^{D\times D}$ shared across the classes $\mathcal{Y}$ given a new dataset $\mathcal{D}^{\oplus}$ based on the Gauss-Newton algorithm leading to:
@@
where $\delta(\cdot)$ is the Dirac delta function. We efficiently compute the updated inverse Hessian using the Woodbury identity as presented in the upcoming Appendix D.
@@ Appendix D Efficient Hessian Inversion via Woodbury Identity
Here, we provide the derivation of the update from Eq. (11), which uses the Woodbury (matrix) identity [55] for efficient inversion of the Hessian during updates. Assume we employed an LA and have the current approximate posterior distribution $q(\boldsymbol{\omega}|\mathcal{D})=\mathcal{N}(\boldsymbol{\omega}|\hat{{\boldsymbol{\mu}}},\hat{{\boldsymbol{\Sigma}}})$. To incorporate the information of the new dataset $\mathcal{D}^{\oplus}$ into the LA’s current covariance $\hat{{\boldsymbol{\Sigma}}}$, we first need to calculate the Hessian of the new negative log posterior $q(\boldsymbol{\omega}|\mathcal{D},\mathcal{D}^{\oplus})$ with respect to $\boldsymbol{\omega}$. The Hessian of the new negative log posterior $-q(\boldsymbol{\omega}|\mathcal{D},\mathcal{D}^{\oplus})$ is given by
${\boldsymbol{H}}=-\nabla^{2}_{\boldsymbol{\omega}}\log q(\boldsymbol{\omega}|\mathcal{D},\mathcal{D}^{\oplus})$ $={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{I}+\sum_{({\boldsymbol{x}},y)\in\mathcal{D}}p_{\boldsymbol{x}}(1-p_{\boldsymbol{x}}){\boldsymbol{h}}_{\boldsymbol{x}}{\boldsymbol{h}}^{T}_{\boldsymbol{x}}}+\sum_{({\boldsymbol{x}},y)\in\mathcal{D}^{\oplus}}p_{\boldsymbol{x}}(1-p_{\boldsymbol{x}}){\boldsymbol{h}}_{\boldsymbol{x}}{\boldsymbol{h}}^{T}_{\boldsymbol{x}},$ (29)
@@
${\boldsymbol{H}}^{-1}=\left(\hat{{\boldsymbol{\Sigma}}}^{-1}+p_{\boldsymbol{x}}(1-p_{\boldsymbol{x}}){\boldsymbol{h}}_{\boldsymbol{x}}{\boldsymbol{h}}_{\boldsymbol{x}}^{T}\right)^{-1}=\hat{{\boldsymbol{\Sigma}}}-\frac{p_{\boldsymbol{x}}(1-p_{\boldsymbol{x}})}{1+{\boldsymbol{h}}_{\boldsymbol{x}}\hat{{\boldsymbol{\Sigma}}}{\boldsymbol{h}}_{\boldsymbol{x}}\cdot p_{\boldsymbol{x}}(1-p_{\boldsymbol{x}})}\hat{{\boldsymbol{\Sigma}}}{\boldsymbol{h}}_{\boldsymbol{x}}{\boldsymbol{h}}_{\boldsymbol{x}}\hat{{\boldsymbol{\Sigma}}},$ (32)
which is the update from Eq. (11). In the multi-class setting, we employ the covariance matrix of Eq. (26) by assuming a Gaussian likelihood leading to the following inverse Hessian:
@@ Appendix E Summary of Datasets
CIFAR [24] consists of 60,000 colored images with a low resolution of $32\times 32$. There is a predetermined split of 50,000 images as training instances and 10,000 images as test instances. One variant of this dataset, CIFAR-10, is a coarse-grained task with ten broad classes such as automobile, dog, airplane, and ship. The classes are mutually exclusive meaning there is no overlap between, e.g., automobiles and trucks. Snacks [33] contains images of 20 different classes of snack foods. The 6,745 available images are split into 3,840 images for training, 955 for validation, and 920 for testing. Each image is 256 pixels wide while its height varies from 256 to 873 pixels. DTD [5] is a texture-focused dataset that consists of 5,640 images divided into 47 different classes. There are 120 images for each class and image sizes range between 300x300 and 640x640. There is a predetermined split into three equally sized datasets for training, validation, and testing. DBPedia [2] is a larger text dataset with a medium class cardinality of 14 different classes. There are 14 different ontology-based classes. It consists of 560,000 training samples and 5,000 test samples. Banking-77 [4] comes with a more complex task of conversational language understanding alongside intent detection. The combination of its high-class cardinality and a small pool of just 10,000 samples makes this data set even more challenging. Clinc-150 [26] includes queries that are out-of-scope, i.e., queries that do not fall into any of the system’s supported intents. It contains 150 in-scope intent classes, each with 100 train, 20 validation, and 30 test instances. Additionally, there are 100 train and validation out-of-scope instances, and 1,000 out-of-scope test instances.
Appendix F Active Learning: Use Case

### v3

@@ 5 Deep Active Learning
Figure 6: Accuracy improvement curves for different datasets showing the accuracy difference between the respective selection strategy and random instance selection.
Improved Batch Selection via Updates: A naive and suboptimal way of using sequential selection strategies for batch selection is to use the top-$b$ scoring instances [21]. Our idea is to overcome the necessity of batch strategies by using the proposed update with sequential strategies as a fast alternative to retraining. Thus, we iteratively select the highest-scoring instance $b$ times and update the DNN between each selection. After acquiring $b$ labels, we retrain the DNN similar to batch selection strategies. An algorithm can be found in Appendix˜F.
The hypothesis is that already well-performing sequential selection strategies [38] can simply be used in a batch setting and that our framework can achieve higher performance compared to selecting the top-$b$ instances. Here, we consider the widely used strategy Margin, which has proven to be effective in several studies [3]. Additionally, we are interested in whether this idea can also replace the diversity component of a batch selection strategy. Therefore, we also evaluate the popular strategy Badge [1] in combination with our updates.
@@
Figure˜6 shows the accuracy improvement curves relative to a random instance selection. The query strategies using our updates outperform the respective top-$b$ selection strategies. Specifically, we see improved performance in early stages when redundancy within a batch plays an important role. Moreover, combining our update with Badge also results in improved accuracy. This indicates that selecting a single instance and updating the DNN leads to a more effective selection than using the $k$-means++ algorithm as proposed in Badge.
Updating in Look-Ahead Strategies: The idea of look-ahead strategies is to select instances that, once labeled and added to the labeled pool, maximize the performance of the model [35]. Unlike uncertainty- or diversity-based approaches, look-ahead strategies select instances based on an optimal criterion: the model’s actual performance. However, they are often neglected in deep AL due to the high computational requirements. One of the biggest bottlenecks in the selection is retraining. DNNs are not well-suited for this due to their long training process. For this reason, we employ our proposed update to make this feasible.
Figure 7: Accuracy improvement over random selection of popular selection strategies compared to our upper baseline approximating optimal batch selection.
Here, we consider a near-optimal strategy with access to ground truth information, including labels and validation datasets. It can be considered as an upper baseline in deep AL research. For the selection, we randomly sample 2000 subsets, each with a size equal to the acquisition size, and assess how their addition to the labeled pool affects the performance. The batch leading to the highest performance gain is selected. While this approach would traditionally require 2000 times of retraining our update enables the efficient use of this strategy.
We also include the recently proposed Typiclust strategy, which demonstrated strong performance [16], especially in early stages of AL. Figure˜7 presents the resulting accuracy improvement compared to random instance selection. Our optimal strategy, using updates rather than full retraining, performs exceptionally well, consistently outperforming all competitors. Based on these results, we can see that the current selection strategies still have much potential for improvement. Interestingly, we can see that Typiclust’s selection in the early stages of AL seems to be close to an optimal selection but declines in effectiveness in later stages.
6 Conclusion
We proposed an efficient second-order update for DNNs in AL using the Gaussian posterior of a last-layer LA. It achieves low computational complexity through a closed-form computation of the required inverse Hessian. An extensive experimental evaluation showed that the proposed update provides an efficient alternative to retraining. Based on this, we introduced a new batch selection framework by sequentially updating the DNN after each label, offering a new perspective on constructing batches without resorting to heuristics such as clustering. Additionally, we realized a look-ahead strategy as a feasible upper baseline approximating optimal batch selection, highlighting the great potential for improvement in current research on batch selection strategies. In future work, we plan to utilize the proposed updates to enhance look-ahead selection strategies [35] in deep AL. As these strategies are based on decision-theoretic principles, they naturally balance explorative and exploitative instance selection, a key challenge in AL.
@@ A.1 MC-based Bayesian Updates
MC-based Bayesian neural networks (BNNs) such as deep ensembles and MC-Dropout draw samples, or hypotheses, from an (approximate) posterior distribution $q(\boldsymbol{\omega}|\mathcal{D})$. To obtain these samples, deep ensembles train multiple randomly initialized deep neural networks (DNNs), while MC-Dropout randomly sets a portion of parameters to zero for multiple inference steps. We refer to [14] for an in-depth explanation of these techniques.
@@
$q(\boldsymbol{\omega}_{m}|\mathcal{D}^{\oplus},\mathcal{D})=\text{Cat}(m|\hat{{\boldsymbol{p}}}^{\text{upd}})$ $\propto q(\boldsymbol{\omega}_{m}|\mathcal{D})\prod_{({\boldsymbol{x}},y)\in\mathcal{D}^{\oplus}}\hskip-8.53581ptp(y|{\boldsymbol{x}},\boldsymbol{\omega}_{m})$ (13)
$=\hat{p}_{m}\prod_{({\boldsymbol{x}},y)\in\mathcal{D}^{\oplus}}\hskip-8.53581pt\left[\mathrm{softmax}(f^{\boldsymbol{\omega}_{m}}({\boldsymbol{x}}))\right]_{y}=\hat{z}_{m},$ (14)
which is a categorical distribution with parameters $\hat{{\boldsymbol{p}}}^{\text{upd}}=(\hat{p}^{\mathrm{upd}}_{1},\ldots,\hat{p}^{\mathrm{upd}}_{M})^{\mathrm{T}}$ that we obtain after normalizing ${\boldsymbol{\hat{z}}}=(\hat{z}_{1},\ldots,\hat{z}_{M})^{\mathrm{T}}$. Intuitively, the importance of each hypothesis is determined by its likelihood of explaining the new dataset $\mathcal{D}^{\oplus}$. This approximation $q(\boldsymbol{\omega}_{m}|\mathcal{D}^{\oplus},\mathcal{D})$ of the posterior distribution allows us to make new predictions by evaluating the predictive distribution from Eq.˜1 accordingly:
@@
which is a weighted average of the sampled hypotheses. Employing the update from Eq.˜14 may lead to catastrophic forgetting. Thus, to control the influence of the new dataset $\mathcal{D}^{\oplus}$, we introduce the hyperparameter $\gamma$:
$q(\boldsymbol{\omega}_{m}|\mathcal{D}^{\oplus},\mathcal{D})$ $\propto\hat{p}_{m}\left(\prod_{({\boldsymbol{x}},y)\in\mathcal{D}^{\oplus}}\hskip-8.53581pt\left[\mathrm{softmax}(f^{\boldsymbol{\omega}_{m}}({\boldsymbol{x}}))\right]_{y}\right)^{\gamma}.$ (17)
@@ A.2 First-Order Bayesian Updates
Our proposed update method from the main text approximates the new posterior $q(\boldsymbol{\omega}|\mathcal{D},\mathcal{D}^{\oplus})$ by applying an optimization step via Gauss-Newton and estimating the new covariance. For comparison, we evaluate an update based on a first-order optimization step, leading to a less complex and faster method. This also allows us to ablate the importance of the Hessian.
[…]
$q(\boldsymbol{\omega}|\mathcal{D},\mathcal{D}^{\oplus})=\mathcal{N}(\boldsymbol{\omega}|\hat{{\boldsymbol{\mu}}}^{\text{upd}},\hat{{\boldsymbol{\Sigma}}}^{\text{upd}})\mathrel{\vbox{\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}q(\boldsymbol{\omega}|\mathcal{D})\prod_{({\boldsymbol{x}},y)\in\mathcal{D}^{\oplus}}p(y|{\boldsymbol{x}},\boldsymbol{\omega}).$ (20)
@@
The updating strategy in Eq.˜21 is equivalent to the continual learning strategy from [34]. There, the updated posterior distribution is defined as
[…]
where the second term is the prior $q(\boldsymbol{\omega}|\mathcal{D})$, penalizing large deviations of $\boldsymbol{\omega}$ from the prior mean $\hat{{\boldsymbol{\mu}}}$. The updated mean is then given by
[…]
[34] use multiple optimization steps via gradient descent to obtain the new mean $\hat{{\boldsymbol{\mu}}}^{\text{upd}}$. Since we defined an update as a single optimization step and we start with $\boldsymbol{\omega}=\hat{{\boldsymbol{\mu}}}$, the regularization term evaluates to zero. Thus, the updated mean is obtained through a gradient step on the new data likelihood, simplifying the optimization to Eq.˜21.
@@ Appendix B Hyperparameter Ablation
For all introduced Bayesian update methods, the hyperparameter $\gamma$ controls the influence of the new dataset $\mathcal{D}^{\oplus}$ on the posterior distribution $q(\boldsymbol{\omega}|\mathcal{D})$. In Sec.˜4, we conducted a small ablation study, similar to a hyperparameter search, to determine an appropriate value for $\gamma$. We intentionally did not search a optimal value of $\gamma$ for every dataset since extensive hyperparameter search for update methods is impractical in an online setting [9]. Using the same setup as in Sec.˜4, we consider the CIFAR-10 and DBPedia datasets. Our baseline DNN is trained on a randomly sampled initial dataset $\mathcal{D}$ of 50 instances. We then update and retrain the DNN with varying sizes of the new dataset $|\mathcal{D}^{\oplus}|\in\{1,\dots,10\}$.
@@ Appendix C Multi-class Last-Layer Update
[39] presented a more efficient approximation in which the class-wise parameter vectors are arranged column-wisely as a matrix. Their joint distribution is then modeled through a matrix normal distribution:
@@
[26] presented an even faster approximation and showed its effectiveness in combination with an LA for supervised learning. Therefore, we employ this approximation in the multi-class setting. The idea is to determine an upper-bound covariance matrix shared by all class-wise parameter vectors in their respective multivariate normal distribution, which is then defined for class $y\in\mathcal{Y}$ as:
@@
Analog to the updates for binary classification in Eq.˜10, we implement the updates of the mean parameter vector for class $y\in\mathcal{Y}$ and the covariance matrix $\hat{{\boldsymbol{\Sigma}}}\in\mathbb{R}^{D\times D}$ shared across the classes $\mathcal{Y}$ given a new dataset $\mathcal{D}^{\oplus}$ based on the Gauss-Newton algorithm leading to:
@@
where $\delta(\cdot)$ is the Dirac delta function. We efficiently compute the updated inverse Hessian using the Woodbury identity as presented in the upcoming Appendix˜D.
@@ Appendix D Efficient Hessian Inversion via Woodbury Identity
[…]
${\boldsymbol{H}}$ $=-\nabla^{2}_{\boldsymbol{\omega}}\log q(\boldsymbol{\omega}|\mathcal{D},\mathcal{D}^{\oplus})$ (31)
$={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{I}+\sum_{({\boldsymbol{x}},y)\in\mathcal{D}}p_{\boldsymbol{x}}(1-p_{\boldsymbol{x}}){\boldsymbol{h}}_{\boldsymbol{x}}{\boldsymbol{h}}^{T}_{\boldsymbol{x}}}+\sum_{({\boldsymbol{x}},y)\in\mathcal{D}^{\oplus}}p_{\boldsymbol{x}}(1-p_{\boldsymbol{x}}){\boldsymbol{h}}_{\boldsymbol{x}}{\boldsymbol{h}}^{T}_{\boldsymbol{x}},$ (32)
@@
${\boldsymbol{H}}^{-1}$ $=\left(\hat{{\boldsymbol{\Sigma}}}^{-1}+p_{\boldsymbol{x}}(1-p_{\boldsymbol{x}}){\boldsymbol{h}}_{\boldsymbol{x}}{\boldsymbol{h}}_{\boldsymbol{x}}^{T}\right)^{-1}$ (35)
$=\hat{{\boldsymbol{\Sigma}}}-\frac{p_{\boldsymbol{x}}(1-p_{\boldsymbol{x}})}{1+{\boldsymbol{h}}_{\boldsymbol{x}}\hat{{\boldsymbol{\Sigma}}}{\boldsymbol{h}}_{\boldsymbol{x}}\cdot p_{\boldsymbol{x}}(1-p_{\boldsymbol{x}})}\hat{{\boldsymbol{\Sigma}}}{\boldsymbol{h}}_{\boldsymbol{x}}{\boldsymbol{h}}_{\boldsymbol{x}}\hat{{\boldsymbol{\Sigma}}}$, (36)
which is the update from Eq.˜11. In the multi-class setting, we employ the covariance matrix of Eq.˜28 by assuming a Gaussian likelihood leading to the following inverse Hessian:
@@ Appendix E Summary of Datasets
CIFAR [23] consists of 60,000 colored images with a low resolution of $32\times 32$. There is a predetermined split of 50,000 images as training instances and 10,000 images as test instances. One variant of this dataset, CIFAR-10, is a coarse-grained task with ten broad classes such as automobile, dog, airplane, and ship. The classes are mutually exclusive meaning there is no overlap between, e.g., automobiles and trucks. Snacks [29] contains images of 20 different classes of snack foods. The 6,745 available images are split into 3,840 images for training, 955 for validation, and 920 for testing. Each image is 256 pixels wide while its height varies from 256 to 873 pixels. DTD [5] is a texture-focused dataset that consists of 5,640 images divided into 47 different classes. There are 120 images for each class and image sizes range between 300x300 and 640x640. There is a predetermined split into three equally sized datasets for training, validation, and testing. DBPedia [2] is a larger text dataset with a medium class cardinality of 14 different classes. There are 14 different ontology-based classes. It consists of 560,000 training instances and 5,000 test instances. Banking-77 [4] comes with a more complex task of conversational language understanding alongside intent detection. The combination of its high-class cardinality and a small pool of just 10,000 instances makes this data set even more challenging. Clinc-150 [25] includes queries that are out-of-scope, i.e., queries that do not fall into any of the system’s supported intents. It contains 150 in-scope intent classes, each with 100 train, 20 validation, and 30 test instances. Additionally, there are 100 train and validation out-of-scope instances, and 1,000 out-of-scope test instances.
Appendix F Active Learning: Updates for Batch Selection and Look-Ahead